# Opportunities and Challenges of Large Language Models for Low-Resource Languages in Humanities Research

### v3

@@ 1.1.1 Importance and Endangerment of Low-Resource Languages in the Global Linguistic Ecology
However, the global linguistic landscape is presently undergoing an extraordinary crisis, with low-resource languages among the most threatened. The swift vanishing of these languages is of serious concern, highlighted by concerning data and studies. It is estimated, for example, that around 40% of the world’s 7,000 languages face extinction, with numerous low-resource languages having fewer than 1,000 speakers [133]. This decline is caused by several factors, such as the widespread effects of globalization, the challenges of urbanization, and the prevalence of dominant languages. These elements often marginalize low-resource languages, resulting in their gradual decline and eventual disappearance.
@@ 1.1.2 Demand for Low-Resource Languages in Humanities Research and the Insufficiencies in Existing Studies
Languages with limited resources are crucial for advancing research in the humanities, especially in fields like anthropology, history, literature, and linguistics. These languages provide distinctive perspectives on cultures, societies, and intellectual traditions that are less frequently studied, thus broadening our appreciation of human diversity. For example, anthropologists frequently use indigenous languages to reveal hidden facets of cultural practices and social organizations not captured by dominant languages. Likewise, historians and literary scholars gain access to primary sources and oral histories maintained in low-resource languages, offering a more detailed and thorough understanding of historical events and literary traditions. In the field of linguistics, analyzing low-resource languages aids the creation of theories and models applicable to a wider variety of human languages, thereby deepening our knowledge of language universals and diversity [49].
The critical significance of low-resource languages notwithstanding, prevailing research methodologies encounter appreciable challenges when engaging with these languages. A principal limitation is the paucity of textual data. Numerous low-resource languages lack extensive written corpora, thereby complicating the execution of empirical studies and the development of computational models. Furthermore, the absence of specialized computational tools specifically adapted to these languages aggravates the issue [83]. Existing tools and resources are frequently designed for high-resource languages, thereby proving ineffective or inefficient for application to low-resource languages. As a brief empirical note from our working list, coverage remains uneven across LLM systems. For example, GPT5 supports Manchu whereas NLLB does not, while NLLB supports Istro-Venetian but GPT5 is not accurate. In short, each system covers some different languages, yet overall support for low-resource, including historical and regional, varieties remains insufficient for dependable humanities workflows. This technological disparity is further heightened by the inadequate availability of essential linguistic resources such as dictionaries, grammars, and annotated corpora, which are imperative for both descriptive and computational linguistic research.
@@ 1.2.1 Breakthroughs in Language Processing with LLMs
The emergence of LLMs marks a new phase in natural language processing (NLP), significantly altering the field with their remarkable capabilities and adaptability. Examples of LLMs, such as GPT-4 [1] and LLaMA [32], are based on the transformer architecture. This architecture uses self-attention mechanisms to efficiently process and generate text with impressive fluency and contextual awareness. By supporting parallel processing of sequences, transformers overcome the challenges faced by earlier architectures like Recurrent Neural Networks (RNNs) [22] and Long Short-Term Memory (LSTM) networks [47], which struggled with handling long-range dependencies and parallelization.
The advancements made by LLMs across several NLP tasks are remarkable. These models have set new benchmarks in areas such as machine translation, text generation, sentiment analysis, and language comprehension. For example, the BERT model [29], which originates from the transformer architecture, has greatly improved the effectiveness and scalability of NLP systems through self-supervised pretraining. In recent developments, generative models like the GPT series [113, 16] have demonstrated their zero-shot and few-shot in-context learning capabilities, allowing them to tackle various tasks without the need for task-specific fine-tuning. This flexibility highlights the extensive datasets used for pretraining these models, which incorporate a vast array of world knowledge and emerging skills.
@@
The incorporation of LLMs into numerous practical applications is capturing growing interest, significantly affecting a variety of areas including education, healthcare, and robotics [80, 155, 91, 25, 85, 77, 82, 117, 24, 153, 149, 86, 87, 89, 134, 57, 88, 58, 130]. Their capacity for human-like comprehension and reasoning is a stepping stone toward Artificial General Intelligence (AGI), fostering societal progress across numerous disciplines[155, 83, 156, 154, 76, 73, 135]. As LLMs advance, they hold vast potential to enrich research in languages with limited resources, providing novel opportunities to preserve and comprehend the extensive cultural and intellectual treasures embedded in these languages.
@@ 2 Linguistic Variation and Dialect Research in Low-Resource Languages
In the field of sociolinguistics, research on under-described languages has traditionally been constrained to specific tasks such as part-of-speech tagging, text classification, and machine translation[92]. However, the emergence of LLMs, as outlined in the foundational framework, has transformed this landscape. By leveraging their computational power, reasoning capabilities, and advanced methodologies, LLMs confront the challenges posed by limited documented linguistic resources, enabling more comprehensive research on linguistic variation and dialects in low-resource languages.
@@ 2.1.1 Existing Methods of Handling Diverse Variants in LLMs
Transfer Learning and Fine-Tuning LLMs utilize transfer learning to adapt knowledge from high-resource languages to low-resource languages[3], effectively bridging linguistic gaps. Fine-tuning on small, domain-specific datasets helps capture regional and dialectal nuances. Techniques such as Low-Rank Adaptation (LoRA) and QLoRA further enhance fine-tuning efficiency, making this approach computationally accessible for low-resource settings.
Prompt Engineering and In-Context Learning Prompt engineering is crucial for tailoring LLMs outputs to specific dialectal contexts. Minor modifications to the prompts, such as the adjustment of word order or structure, can significantly influence the model output[107]. In-context learning further supports low-resource scenarios by enabling LLMs to adapt without requiring extensive retraining, leveraging contextual examples to generate more accurate responses.
Retrieval-Augmented Generation RAG enhances the performance of LLMs by integrating retrieval mechanisms that fetch relevant contextual data, which is especially valuable for underrepresented languages. It is particularly valuable for underrepresented languages, where contextual augmentation can compensate for limited data availability[125]. In addition to leveraging retrieval data to enhance training datasets, there are advanced RAG approaches that dynamically select the appropriate dialect context from external knowledge bases or corpora to enhance generation accuracy.
@@ 2.1.2 Dialect Speech Recognition in Low-Resource Contexts
Speech recognition systems are indispensable for addressing dialect recognition in low-resource settings, where annotated datasets are scarce. Advanced methods such as the state-of-the-art Conformer Transducer (ConfT) model, combined with Model-Agnostic Meta-Learning (MAML), have demonstrated substantial improvements in dialect-specific tasks. These methods optimize parameters for rapid fine-tuning with minimal data, reducing word error rates by up to 12% for low-resource accents, as evidenced in recent studies[118] [35]. Building on advancements in speech-to-text and text-to-speech models, LLMs can further enhance dialect recognition by integrating phonetic and acoustic features into language understanding pipelines. By leveraging dialect-specific lexicons and fine-tuned language models, these systems achieve improved transcription accuracy, making them vital tools for the preservation and accessibility of linguistic diversity.
@@ 2.1.3 Challenges in Dialect Recognition and Modeling
  • Biases and Sampling Limitations: Over-representation[54] of high-resource languages in training data skews model output, undermining efforts to build inclusive systems capable of capturing the full spectrum of linguistic variation.
@@ 2.2.2 Spatiotemporal variant processing in low-resource languages
Cross-lingual Transfer Learning Cross-lingual transfer learning, as referenced in [19], is a widely used approach that leverages shared features from high-resource languages to aid in the processing of low-resource languages. This method helps capture fundamental linguistic features and can be generalized to accommodate different spatiotemporal variations.
Removal of Explicit Language Tag Embeddings By removing explicit language classification encodings for words, this approach allows different languages to share feature representations[23]. It enables the model to benefit from data across multiple languages, thereby enhancing its broader linguistic understanding.
Multimodal Architectures and Representation Learning Enhancing LLMs’ ability to process low-resource languages through multimodal learning is an effective strategy[95]. For example, combining visual and textual features can improve the model’s understanding ability of spatiotemporal variations and cultural context differences.
Sentence Augmentation Techniques Some studies focus on sentence augmentation techniques to handle low-resource languages, such as Kazakh[13]. By generating more diverse sentences and expressions using LLMs, these methods address regional language variations. They enable the generation of more natural low-resource language sentences, expanding the data coverage and improving the model’s generalization capabilities for new language inputs.
@@ 2.3 Opportunities and Challenges
Automated modeling provide linguistic scholar transformational opportunities with advancing research on low-resource languages like Quechua by enabling in-depth analysis of unique linguistic features and dialectal variability. For instance, it provides tools for interpreting sociolinguistic contexts, such as the placeholder ”na,” which functions both as a hesitation marker and as a substitute for omitted words, offering insights into bilingual code-switching and linguistic structure. Automated approaches also support the exploration of dialectal variability and linguistic borrowing[90]. These contributions not only facilitate the preservation of linguistic diversity, but also enhance understanding of bilingual interactions and address the complexities inherent in dialectal variation.
@@ 3 Applications of Low-Resource Languages in Historical Research
Low-resource languages often carry rich historical and cultural information but are now at risk of gradual extinction. Protecting low-resource languages is not only a critical topic in linguistic research but also an essential aspect of historical studies. These languages have become “low-resource” due to a combination of factors. Over thousands of years, wars, colonialism, and political upheaval have forced certain language communities into displacement, leading to interruptions in language transmission [119, 122, 123, 120, 121]. Industrialization and modernization have transformed traditional agricultural societies, sparking intense social change [142]. Many people have migrated from rural areas to cities, abandoning their original languages in favor of more widely spoken ones. Under the influence of globalization, dominant cultures and languages often overshadow and displace weaker ones.
The importance of protecting low-resource languages within historical research cannot be underestimated. By preserving and studying these languages, we gain a deeper understanding of humanity’s past, enrich historical scholarship, and promote cultural diversity and social inclusivity. With the advancement of artificial intelligence and natural language processing technologies, efforts to protect low-resource languages have gained new momentum [84, 80, 81, 134]. Large language models and computer vision technologies can be used to identify, translate, and interpret low-resource language literature, providing powerful tools for historical research [127]. The application of these technologies not only improves research efficiency but also expands the depth and scope of studies, allowing a broader range of low-resource language literature to be fully utilized. This section will delve into the application of LLMs in the interpretation of historical literature in low-resource languages.
@@ 3.1 Interpretation of Historical Literature
There has already been some work using AI to translate and decode low-resource languages, such as the deciphering and restoration of the Dead Sea Scrolls [112, 30, 31, 69], the recognition of ancient Egyptian hieroglyphs [12], text restoration of ancient Greek inscriptions [10], the decoding of oracle bone script [48], and the unified Visual-Linguistic understanding of oracle bone scripts[61]. These early efforts, however, did not yet make use of LLMs. LLMs can play a pivotal role in addressing these challenges. Firstly, to combat the scarcity of literature, LLMs coupled with advanced computer vision techniques can assist in digitizing and reconstructing fragmented or damaged ancient texts. This not only preserves existing documents but also consolidates them into accessible digital archives for researchers worldwide. Secondly, regarding linguistic complexity, LLMs can be fine-tuned on available data, even if limited, to learn the unique grammatical structures and vocabulary of low-resource languages. Lastly, to mitigate the scarcity of language experts, LLMs can serve as virtual assistants, providing preliminary translations and analyses. This empowers a broader range of scholars and enthusiasts to engage with these languages, thereby expanding the pool of individuals who can contribute to their study and preservation.
@@ 3.2 Data Mining and Knowledge Reconstruction in Historical Contexts
LLMs have demonstrated remarkable potential in the mining of historical data and the restoration of knowledge, particularly concerning low-resource languages [127]. These advanced models can learn the unique grammatical structures and vocabularies of these languages from limited datasets. By doing so, they can fill information gaps in historical documents and unearth unsystematized historical knowledge and informal records that have long been inaccessible to researchers.
@@
LLMs can process and analyze these informal records, extracting historical information that might otherwise remain hidden. By identifying patterns and drawing connections across disparate documents, LLMs provide new research perspectives and enable historians to construct more nuanced narratives [80]. This capability is particularly crucial for low-resource languages, where such informal records may be the primary sources of historical data.
@@ 3.3 Challenges and Opportunities
The application of LLMs to the historical study of low-resource languages has shown great potential, but still faces many technical difficulties in practice. Among them, the limitations of optical character recognition (OCR) technology, a key technology for digitizing ancient documents, are particularly prominent. The standard processing flow of OCR includes image preprocessing, text region localization, feature extraction, character recognition, and post-processing [11]. In these links, the accurate positioning of text region is crucial, which mainly involves two tasks of text detection and character segmentation.
The special characteristics of ancient documents make these two tasks more difficult to accomplish. Many documents are presented in handwritten or non-standardized fonts, and are often difficult to recognize due to blurring, fading, or physical damage [99]. Such conditions place high demands on OCR systems, making it difficult to effectively recognize areas containing text during text detection [101]. Similarly, some ancient documents (e.g., cursive scripts) have continuous writing and characters stacked on top of each other, which poses a challenge for character segmentation. These scenarios require the system to have higher resolution capability and accurate segmentation algorithms to avoid character misjudgment or character concatenation phenomenon.
Last but not least, there are many limitations in automatically converting ancient characters to modern languages. Low-resource languages often lack sufficient text corpora to provide high-quality training data for LLMs [92], which results in lower accuracy rates for translation or transcription of ancient characters. To further complicate matters, these languages tend to be morphologically variable and complex, and lack standardized writing rules. The morphology of a word may be varied and diverse depending on the context, tense, or syntactic roles, which poses a great problem for models to recognize and translate ancient texts. Finally, most of the current LLMs are primarily trained on large-scale datasets of modern languages, so researchers need to develop new approaches or adapt existing technical frameworks to address the particular challenges of low-resource scripts.

### v4

@@ 1.1.1 Importance and Endangerment of Low-Resource Languages in the Global Linguistic Ecology
However, the global linguistic landscape is presently undergoing an extraordinary crisis, with low-resource languages among the most threatened. The swift vanishing of these languages is of serious concern, highlighted by concerning data and studies. It is estimated, for example, that around 40% of the world’s 7,000 languages face extinction, with numerous low-resource languages having fewer than 1,000 speakers [vasconcelos2024disappearing]. This decline is caused by several factors, such as the widespread effects of globalization, the challenges of urbanization, and the prevalence of dominant languages. These elements often marginalize low-resource languages, resulting in their gradual decline and eventual disappearance.
@@ 1.1.2 Demand for Low-Resource Languages in Humanities Research and the Insufficiencies in Existing Studies
Languages with limited resources are crucial for advancing research in the humanities, especially in fields like anthropology, history, literature, and linguistics. These languages provide distinctive perspectives on cultures, societies, and intellectual traditions that are less frequently studied, thus broadening our appreciation of human diversity. For example, anthropologists frequently use indigenous languages to reveal hidden facets of cultural practices and social organizations not captured by dominant languages. Likewise, historians and literary scholars gain access to primary sources and oral histories maintained in low-resource languages, offering a more detailed and thorough understanding of historical events and literary traditions. In the field of linguistics, analyzing low-resource languages aids the creation of theories and models applicable to a wider variety of human languages, thereby deepening our knowledge of language universals and diversity [gwerevende2023safeguarding].
The critical significance of low-resource languages notwithstanding, prevailing research methodologies encounter appreciable challenges when engaging with these languages. A principal limitation is the paucity of textual data. Numerous low-resource languages lack extensive written corpora, thereby complicating the execution of empirical studies and the development of computational models. Furthermore, the absence of specialized computational tools specifically adapted to these languages aggravates the issue [liu2023transformation]. Existing tools and resources are frequently designed for high-resource languages, thereby proving ineffective or inefficient for application to low-resource languages. As a brief empirical note from our working list, coverage remains uneven across LLM systems. For example, GPT5 supports Manchu whereas NLLB does not, while NLLB supports Istro-Venetian but GPT5 is not accurate. In short, each system covers some different languages, yet overall support for low-resource, including historical and regional, varieties remains insufficient for dependable humanities workflows. This technological disparity is further heightened by the inadequate availability of essential linguistic resources such as dictionaries, grammars, and annotated corpora, which are imperative for both descriptive and computational linguistic research.
@@ 1.2.1 Breakthroughs in Language Processing with LLMs
The emergence of LLMs marks a new phase in natural language processing (NLP), significantly altering the field with their remarkable capabilities and adaptability. Examples of LLMs, such as GPT-4 [achiam2023gpt] and LLaMA [dubey2024llama], are based on the transformer architecture. This architecture uses self-attention mechanisms to efficiently process and generate text with impressive fluency and contextual awareness. By supporting parallel processing of sequences, transformers overcome the challenges faced by earlier architectures like Recurrent Neural Networks (RNNs) [cho2014learning] and Long Short-Term Memory (LSTM) networks [graves2012long], which struggled with handling long-range dependencies and parallelization.
The advancements made by LLMs across several NLP tasks are remarkable. These models have set new benchmarks in areas such as machine translation, text generation, sentiment analysis, and language comprehension. For example, the BERT model [devlin2018bert], which originates from the transformer architecture, has greatly improved the effectiveness and scalability of NLP systems through self-supervised pretraining. In recent developments, generative models like the GPT series [radford2019language, brown2020language] have demonstrated their zero-shot and few-shot in-context learning capabilities, allowing them to tackle various tasks without the need for task-specific fine-tuning. This flexibility highlights the extensive datasets used for pretraining these models, which incorporate a vast array of world knowledge and emerging skills.
@@
The incorporation of LLMs into numerous practical applications is capturing growing interest, significantly affecting a variety of areas including education, healthcare, and robotics [liu2023summary, zhao2023brain, ma2024iterative, dai2023chataug, liu2023radiology, liao2023differentiating, liu2023context, rezayi2022clinicalradiobert, dai2023ad, zhao2023ophtha, zhang2024generalist, liu2023radonc, liu2024fine, lyu2024gp, wang2024comprehensive, huang2024position, liu2024surviving, huang2024trustllm, tian2024assessing]. Their capacity for human-like comprehension and reasoning is a stepping stone toward Artificial General Intelligence (AGI), fostering societal progress across numerous disciplines[zhao2023brain, liu2023transformation, zhenyuan2024analyzing, zhao2024revolutionizing, li2024large, lee2023multimodality, wang2024legal]. As LLMs advance, they hold vast potential to enrich research in languages with limited resources, providing novel opportunities to preserve and comprehend the extensive cultural and intellectual treasures embedded in these languages.
@@ 2 Linguistic Variation and Dialect Research in Low-Resource Languages
In the field of sociolinguistics, research on under-described languages has traditionally been constrained to specific tasks such as part-of-speech tagging, text classification, and machine translation[magueresse2020lowresourcelanguagesreviewpast]. However, the emergence of LLMs, as outlined in the foundational framework, has transformed this landscape. By leveraging their computational power, reasoning capabilities, and advanced methodologies, LLMs confront the challenges posed by limited documented linguistic resources, enabling more comprehensive research on linguistic variation and dialects in low-resource languages.
@@ 2.1.1 Existing Methods of Handling Diverse Variants in LLMs
Transfer Learning and Fine-Tuning LLMs utilize transfer learning to adapt knowledge from high-resource languages to low-resource languages[alam2024llms], effectively bridging linguistic gaps. Fine-tuning on small, domain-specific datasets helps capture regional and dialectal nuances. Techniques such as Low-Rank Adaptation (LoRA) and QLoRA further enhance fine-tuning efficiency, making this approach computationally accessible for low-resource settings.
Prompt Engineering and In-Context Learning Prompt engineering is crucial for tailoring LLMs outputs to specific dialectal contexts. Minor modifications to the prompts, such as the adjustment of word order or structure, can significantly influence the model output[parida2024pretrain]. In-context learning further supports low-resource scenarios by enabling LLMs to adapt without requiring extensive retraining, leveraging contextual examples to generate more accurate responses.
Retrieval-Augmented Generation RAG enhances the performance of LLMs by integrating retrieval mechanisms that fetch relevant contextual data, which is especially valuable for underrepresented languages. It is particularly valuable for underrepresented languages, where contextual augmentation can compensate for limited data availability[seo2024retrievalaugmenteddataaugmentationlowresource]. In addition to leveraging retrieval data to enhance training datasets, there are advanced RAG approaches that dynamically select the appropriate dialect context from external knowledge bases or corpora to enhance generation accuracy.
@@ 2.1.2 Dialect Speech Recognition in Low-Resource Contexts
Speech recognition systems are indispensable for addressing dialect recognition in low-resource settings, where annotated datasets are scarce. Advanced methods such as the state-of-the-art Conformer Transducer (ConfT) model, combined with Model-Agnostic Meta-Learning (MAML), have demonstrated substantial improvements in dialect-specific tasks. These methods optimize parameters for rapid fine-tuning with minimal data, reducing word error rates by up to 12% for low-resource accents, as evidenced in recent studies[romanenko2020robust] [10485786]. Building on advancements in speech-to-text and text-to-speech models, LLMs can further enhance dialect recognition by integrating phonetic and acoustic features into language understanding pipelines. By leveraging dialect-specific lexicons and fine-tuned language models, these systems achieve improved transcription accuracy, making them vital tools for the preservation and accessibility of linguistic diversity.
@@ 2.1.3 Challenges in Dialect Recognition and Modeling
  • Biases and Sampling Limitations: Over-representation[terhoeve2022highresourcemethodologicalbiaslowresource] of high-resource languages in training data skews model output, undermining efforts to build inclusive systems capable of capturing the full spectrum of linguistic variation.
@@ 2.2.2 Spatiotemporal variant processing in low-resource languages
Cross-lingual Transfer Learning Cross-lingual transfer learning, as referenced in [cekinel2024cross], is a widely used approach that leverages shared features from high-resource languages to aid in the processing of low-resource languages. This method helps capture fundamental linguistic features and can be generalized to accommodate different spatiotemporal variations.
Removal of Explicit Language Tag Embeddings By removing explicit language classification encodings for words, this approach allows different languages to share feature representations[conneau2019unsupervised]. It enables the model to benefit from data across multiple languages, thereby enhancing its broader linguistic understanding.
Multimodal Architectures and Representation Learning Enhancing LLMs’ ability to process low-resource languages through multimodal learning is an effective strategy[mani2023large]. For example, combining visual and textual features can improve the model’s understanding ability of spatiotemporal variations and cultural context differences.
Sentence Augmentation Techniques Some studies focus on sentence augmentation techniques to handle low-resource languages, such as Kazakh[bimagambetova2023evaluating]. By generating more diverse sentences and expressions using LLMs, these methods address regional language variations. They enable the generation of more natural low-resource language sentences, expanding the data coverage and improving the model’s generalization capabilities for new language inputs.
@@ 2.3 Opportunities and Challenges
Automated modeling provide linguistic scholar transformational opportunities with advancing research on low-resource languages like Quechua by enabling in-depth analysis of unique linguistic features and dialectal variability. For instance, it provides tools for interpreting sociolinguistic contexts, such as the placeholder ”na,” which functions both as a hesitation marker and as a substitute for omitted words, offering insights into bilingual code-switching and linguistic structure. Automated approaches also support the exploration of dialectal variability and linguistic borrowing[languages9030082]. These contributions not only facilitate the preservation of linguistic diversity, but also enhance understanding of bilingual interactions and address the complexities inherent in dialectal variation.
@@ 3 Applications of Low-Resource Languages in Historical Research
Low-resource languages often carry rich historical and cultural information but are now at risk of gradual extinction. Protecting low-resource languages is not only a critical topic in linguistic research but also an essential aspect of historical studies. These languages have become “low-resource” due to a combination of factors. Over thousands of years, wars, colonialism, and political upheaval have forced certain language communities into displacement, leading to interruptions in language transmission [saunt1999new, saunt2017our, saunt2020unworthy, saunt2005black, saunt2016age]. Industrialization and modernization have transformed traditional agricultural societies, sparking intense social change [yang2021medicated]. Many people have migrated from rural areas to cities, abandoning their original languages in favor of more widely spoken ones. Under the influence of globalization, dominant cultures and languages often overshadow and displace weaker ones.
The importance of protecting low-resource languages within historical research cannot be underestimated. By preserving and studying these languages, we gain a deeper understanding of humanity’s past, enrich historical scholarship, and promote cultural diversity and social inclusivity. With the advancement of artificial intelligence and natural language processing technologies, efforts to protect low-resource languages have gained new momentum [liu2023, liu2023summary, liu2024understanding, wang2024comprehensive]. Large language models and computer vision technologies can be used to identify, translate, and interpret low-resource language literature, providing powerful tools for historical research [sommerschield2023machine]. The application of these technologies not only improves research efficiency but also expands the depth and scope of studies, allowing a broader range of low-resource language literature to be fully utilized. This section will delve into the application of LLMs in the interpretation of historical literature in low-resource languages.
@@ 3.1 Interpretation of Historical Literature
There has already been some work using AI to translate and decode low-resource languages, such as the deciphering and restoration of the Dead Sea Scrolls [popovic2021artificial, dhali2017digital, dhali2019binet, koopmans2024performance], the recognition of ancient Egyptian hieroglyphs [barucci2022ancient], text restoration of ancient Greek inscriptions [assael2022restoring], the decoding of oracle bone script [guan2024deciphering], and the unified Visual-Linguistic understanding of oracle bone scripts[jiang2024oraclesageunifiedvisuallinguisticunderstanding]. These early efforts, however, did not yet make use of LLMs. LLMs can play a pivotal role in addressing these challenges. Firstly, to combat the scarcity of literature, LLMs coupled with advanced computer vision techniques can assist in digitizing and reconstructing fragmented or damaged ancient texts. This not only preserves existing documents but also consolidates them into accessible digital archives for researchers worldwide. Secondly, regarding linguistic complexity, LLMs can be fine-tuned on available data, even if limited, to learn the unique grammatical structures and vocabulary of low-resource languages. Lastly, to mitigate the scarcity of language experts, LLMs can serve as virtual assistants, providing preliminary translations and analyses. This empowers a broader range of scholars and enthusiasts to engage with these languages, thereby expanding the pool of individuals who can contribute to their study and preservation.
@@ 3.2 Data Mining and Knowledge Reconstruction in Historical Contexts
LLMs have demonstrated remarkable potential in the mining of historical data and the restoration of knowledge, particularly concerning low-resource languages [sommerschield2023machine]. These advanced models can learn the unique grammatical structures and vocabularies of these languages from limited datasets. By doing so, they can fill information gaps in historical documents and unearth unsystematized historical knowledge and informal records that have long been inaccessible to researchers.
@@
LLMs can process and analyze these informal records, extracting historical information that might otherwise remain hidden. By identifying patterns and drawing connections across disparate documents, LLMs provide new research perspectives and enable historians to construct more nuanced narratives [liu2023summary]. This capability is particularly crucial for low-resource languages, where such informal records may be the primary sources of historical data.
@@ 3.3 Challenges and Opportunities
The application of LLMs to the historical study of low-resource languages has shown great potential, but still faces many technical difficulties in practice. Among them, the limitations of optical character recognition (OCR) technology, a key technology for digitizing ancient documents, are particularly prominent. The standard processing flow of OCR includes image preprocessing, text region localization, feature extraction, character recognition, and post-processing [avyodri2022optical]. In these links, the accurate positioning of text region is crucial, which mainly involves two tasks of text detection and character segmentation.
The special characteristics of ancient documents make these two tasks more difficult to accomplish. Many documents are presented in handwritten or non-standardized fonts, and are often difficult to recognize due to blurring, fading, or physical damage [memon2020handwritten]. Such conditions place high demands on OCR systems, making it difficult to effectively recognize areas containing text during text detection [nguyen2021survey]. Similarly, some ancient documents (e.g., cursive scripts) have continuous writing and characters stacked on top of each other, which poses a challenge for character segmentation. These scenarios require the system to have higher resolution capability and accurate segmentation algorithms to avoid character misjudgment or character concatenation phenomenon.
Last but not least, there are many limitations in automatically converting ancient characters to modern languages. Low-resource languages often lack sufficient text corpora to provide high-quality training data for LLMs [magueresse2020lowresourcelanguagesreviewpast], which results in lower accuracy rates for translation or transcription of ancient characters. To further complicate matters, these languages tend to be morphologically variable and complex, and lack standardized writing rules. The morphology of a word may be varied and diverse depending on the context, tense, or syntactic roles, which poses a great problem for models to recognize and translate ancient texts. Finally, most of the current LLMs are primarily trained on large-scale datasets of modern languages, so researchers need to develop new approaches or adapt existing technical frameworks to address the particular challenges of low-resource scripts.